Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-08-22

Observers: David Homan, Joe Lyman
TAT: Morgan Fraser, Giorgos Leloudas
Support team: Morgan Fraser

Calibrations: biases, dome arcs

Conditions: Closed at start of night, heavy cloud cover. Quite high winds, ~10m/s. Passing clouds for the short duration dome was opened. Close for 2nd half of night due to thick clouds.

UT    Target            Type        Comment
==============================================================================================
02:30                               Dome opened under still poor conditions, switch to SOFI.
02:35    AT2017gah      followup    JHK - very marginal detection in J exposures only.
02:45    AT2017gah      followup    Started BG OB - no guide star and conditions very poor. Didn't start spectrum, instead repeated JHK for AT2017gah with extended OBs - clouds hit and nothing in exposures.
03:20    AT2017gah      followup    JHK again. Conditions improved a bit, good detection in individual J frames, seems to be there in HK
03:50    AT2016iks      followup    JHK. K band in particular was affected by incoming clouds
05:15                               Nothing visible on all-sky cam. Wait out clouds.
05:40                               Dome closed.
09:30                               Dome opened.
09:35    AT2017gax      followup    JHK. JH detections in frames by eye, just about there in K
10:00    AT2017fzw      followup    JHK. J detections in frames by eye, just about there in HK

Comments